# FedReFT: Federated Representation Fine-Tuning with All-But-Me Aggregation

### v1

@@ 3.3 All-But-Me (ABM) Aggregation
To solve this optimization efficiently, we employ Weiszfeld’s algorithm Weiszfeld (1937), an iterative method known to converge under mild conditions. Details of the algorithm are provided in Appendix E. By avoiding direct averaging and incorporating semantically meaningful low-rank intervention updates through robust aggregation, ABM enables each client to benefit from the knowledge of others without sacrificing local personalization. This approach enhances stability and generalization across non-i.i.d. and task-heterogeneous FL environments.
@@
Table 3: Federated fine-tuning performance of LlaMa-3.2 3B across five commonsense reasoning tasks with MT experimental setup where clients train on heterogeneous task mixtures to promote generalizable representations. *Performance results of all baseline methods and the experimental setup are taken from Singhal et al. (2025).
Method R # TP(M) $\downarrow$ TP(%) BoolQ PIQA SIQA HellaS. WinoG Avg $\uparrow$
FLoRA* 32 243.15 7.58 65.05 82.81 74.67 81.84 76.01 78.83
FedIT* 32 48.63 1.51 62.99 81.50 73.13 76.83 71.51 75.74
FFA-LoRA* 32 24.31 0.76 62.87 80.03 68.53 70.02 65.56 71.11
Fed-SB * 120 2.83 0.0884 64.86 81.66 74.87 81.67 75.22 75.66
FedReFT (Ours) 4 1.38 0.0428 63.09 82.10 72.36 90.27 69.22 75.41
8 2.75 0.0857 64.01 81.18 72.11 90.71 71.01 75.66
16 6.194 0.1927 63.42 81.61 73.64 91.23 71.35 76.05
32 11.01 0.3427 64.53 81.34 73.39 91.51 71.32 76.22
FedReFT (tie $\phi$, Ours) 4 0.688 0.0214 49.94 81.23 72.72 89.84 68.43 72.43
8 1.38 0.0428 57.15 81.22 72.77 90.56 68.50 74.04
Table 4: Performance comparison across arithmetic reasoning tasks with the Distict Task and Mixed Task setup.
@@
Models AQuA GSM8K SVAMP Avg $\uparrow$ AQuA GSM8K SVAMP Avg $\uparrow$
LLaMa 7B 25.59 25.47 49.80 33.62 22.83 14.33 27.10 21.42
LLaMa-2 7B 29.53 32.45 57.3 39.76 21.65 20.39 31.5 24.51
LLaMa-3 8B 34.64 48.98 73.60 52.41 31.89 48.90 70.04 50.48
@@ 4 Experimental Validation
To evaluate FedReFT, we conduct extensive experiments on three different NLP benchmarks covering over 12 datasets. Our objective is to present a comprehensive assessment of how this approach performs in various NLP tasks. We experiment with both masked and autoregressive language models, including RoBERTa-large Liu et al. (2019), TinyLlama-1B Community (2023), LLaMA 7B Touvron et al. (2023a), LLaMA-2 7B and 13B Touvron et al. (2023b), LLaMA-3.2B and LLaMA-3 8B AI (2024), across multiple settings and scales. Our comparisons include state-of-the-art baselines, such as LoRAHu et al. (2021b), FedIT Zhang et al. (2024), FFA-LoRASun et al. (2024), FedDPA-LoRALong et al. (2024), FedSA-LoRAGuo et al. (2024), Fed-SB Singhal et al. (2025) and FLoRAWang et al. (2024) focusing on both parameter efficiency and performance trade-offs. We align the experimental setup configurations with the baseline papers to ensure fair comparisons. To optimize memory usage, we load all base language models with torch.bfloat16 precision. All experiments are executed on a single NVIDIA A100-SXM4-80GB GPU, except for LLaMA-2 13B, which is run on a GPUH200x8 141GB system to accommodate the computational demands of large-scale federated fine-tuning. The results are averaged over two runs to report the mean performance.
Hyperparameter Configuration. In the experiments, we determine how many interventions to learn, as well as which input positions and layers to apply them to. We apply interventions at a fixed number of layers $L$, and at prefix ($p$) and suffix ($s$) positions in the input prompt. We narrow the hyperparameter search space for Federated Learning by adopting the configuration used in the centralized ReFTWu et al. (2024b) paper. The appendix B provides a brief overview of the hyperparameter search space. We experiment with whether to share (tie) the intervention parameters $\phi$ across different input positions within the same layer. Given the positions $P=\{1,\ldots,p\}\cup\{n-s+1,\ldots,n\}$, we define the untied and tied variants Wu et al. (2024b):
Task Distribution Rationale.
We design two experimental setups to study how global representations converge under diverse task distributions. In the Mixed-Task (MT) setup, each client trains on a subset of a combined reasoning dataset but is evaluated on a single task, encouraging generalized, transferable representations through ABM aggregation. This reflects collaborative learning across varied yet related tasks. In the Distinct Task (DT) setup, each client trains on a unique reasoning task, enabling personalized fine-tuning while still leveraging global updates. Despite higher task heterogeneity, this setup maintains stable performance as model capacity increases. Both setups show that FedReFT supports effective generalization in MT and robustness in DT.
@@ 4.1 Commonsense Reasoning
We evaluate global representation generation on eight commonsense reasoning tasks using the Commonsense170K dataset inspired by Singhal et al. (2025); Wu et al. (2024b). We use the same hyperparameter of  Singhal et al. (2025) and tune the intervention parameter in the Appendix B.1. This helps us tune important hyperparameters efficiently and also test their robustness across multiple commonsense reasoning tasks.
Datasets. For the first setup (MT design), we split the combined COMMONSENSE170K Hu et al. (2023) commonsense reasoning tasks among clients and use them for fine-tuning. Each client evaluates one of the commonsense reasoning tasks. BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), and WinoGrande Sakaguchi et al. (2021). For the second setup (DT design), each client fine-tunes on only one of these five commonsense reasoning tasks and is evaluated using the same task. All examples are formatted as multiple-choice questions, requiring the model to directly generate the correct answer without providing rationales. We adopt the prompt template from Hu et al. Hu et al. (2023) with minor modifications, including additional string normalization by removing leading and trailing whitespace.
Table 5: Performance comparison across GLUE Tasks on RoBERTa model for $C=3$, FedReFT use rank rank 1. *Performance results of all baseline methods are taken from Guo et al. (2024) and use LoRA rank 8.
Setup Method # TP(M) $\downarrow$ TP(%) MNLI-m SST-2 QNLI QQP Avg $\uparrow$
Standalone FT 355 100 88.8 96.0 93.8 91.5 91.87
LoRA* 1.83 0.515 88.71 95.16 91.16 85.33 89.33
LoReFT 0.053 0.015 89.2 96.2 94.1 88.5 92.0
FL FFA-LoRA* 1.44 0.405 88.83 94.95 91.52 86.71 89.39
FedDPA-LoRA* 2.62 0.737 88.99 95.50 90.74 85.73 89.47
FedSA-LoRA* 1.83 0.551 90.18 96.00 92.13 87.48 90.43
FedReFT (ours) 0.053 0.015 88.86 95.17 94.52 86.57 90.93
@@
Table 6: Performance comparison on arithmetic reasoning tasks for GSM8K on LLaMa-3 8B model with LoRA rank 8, where clients enable consistent evaluation of representation generalization. *Performance results of all baseline methods are taken from Guo et al. (2024).
Method # TP(M) $\downarrow$ TP(%) GSM8K
LoReFT 4.19 0.052 48.33
LoRA* 30.40 0.38 46.23
FedSA-LoRA* 30.40 0.38 46.63
FFA-LoRA* 15.2 0.19 46.32
FedReFT (tie $\phi$) 2.09 0.0261 49.35
FedReFT 4.19 0.0622 49.68
@@ 4.2 Arithmetic Reasoning
Datasets. In the first setting following MT, we split a combined arithmetic reasoning dataset, MATH10K Hu et al. (2023) which includes four arithmetic reasoning tasks with chain-of-thought solutions generated by a language model. Each client reports performance using test set one of three tasks: AQuA Ling et al. (2017), GSM8K Cobbe et al. (2021), and SVAMP Patel et al. (2021). In the second setting following DT, each client is assigned one arithmetic reasoning task for both fine-tuning and evaluation. Both MT and DT setup results are reported in Table 4. In the third setting following Guo et al. (2024); Kuang et al. (2024), we split the dataset GSM8K into three clients under an IID distribution, and the results are shown in Table LABEL:tbl:math_setup_3_baseline. All optimization hyperparameters remain consistent with that setup.
Results. In Table LABEL:tbl:math_setup_3_baseline, FedReFT demonstrates strong performance while using significantly fewer trainable parameters than existing baselines. Notably, FedReFT achieves the highest accuracy among all methods, while FedReFT (tie $\phi$) offers a compelling trade-off between performance and efficiency. These results highlight the scalability and efficiency of our representation-tuning approach. The Distinct Task (DT) setup represents task heterogeneity. Hence, fine-tuning allows clients to learn highly personalized, task-specific representations while benefiting from global aggregation. As a result, the DT setup yields higher performance, as seen in Table 4. In contrast, the MT setup, where clients train on heterogeneous task mixtures to promote global generalizable representations. This blending of tasks during fine-tuning leads to general representation learning but can degrade performance on specific evaluation tasks due to misaligned representation and conflicting task objectives.
@@ 4.3 Natural Language Understanding
We evaluate the effectiveness of FedReFT in learning generalizable representations for Natural Language Understanding (NLU) using the GLUE benchmark Wang et al. (2018). The objective is to fine-tune NLU to learn global representations that capture task-level semantics. By aligning intermediate representations for downstream classification performance. This setup allows us to test whether lightweight intervention tuning can align representations across clients within a single NLU task.
Hyperparameter Tuning. We tune hyperparameters separately for each task, following common practice for PEFT methods Hu et al. (2023) in FL. To reduce the impact of random seed variability, we perform hyperparameter tuning using a fixed seed and report the average performance across that seed and two additional unseen seeds. Details are provided in the Appendix B.2.
Results. Table LABEL:tbl:glue_task depicts that our approach performs strongly across GLUE tasks while using very few trainable parameters. It performs competitive or outperforms other methods, showing that it can learn good representations even in a federated setting. Despite using over 30× fewer parameters than some baselines, it still achieves competitive results, making it both efficient and effective.
@@ 4.4 Ablation Studies
We conduct ablation studies to better understand the effectiveness of FedReFT, specifically examining the role of geometric median-based All-But-Me aggregation. Details of this analysis are provided in Appendix F.
@@ 5 Conclusion
In this work, we addressed a critical gap in the deployment of Representation Fine-Tuning within Federated Learning settings by proposing a novel aggregation strategy tailored to its low-rank, representation-level interventions. While ReFT improves upon traditional PEFT methods like LoRA by operating on semantically rich hidden representations, its application in FL is limited by data heterogeneity, model diversity, and the shortcomings of standard aggregation methods. To address these, we propose All-But-Me aggregation, enabling clients to adapt their local ReFT parameters using a robust average of others’ interventions. FedReFT ensures parameter efficiency and semantic alignment. Extensive experiments under task heterogeneity and different heterogeneous settings show that ABM consistently enhances convergence, generalization, and robustness, making it a practical and effective solution for personalized representation learning in federated systems.

### v2

@@ 3.3 All-But-Me (ABM) Aggregation on Server
Table 1: Comparison of FedAvg, Mean_ABM, and GeoMedian_ABM across commonsense, arithmetic, and GLUE tasks. Geometric median-based ABM aggregation achieves better accuracy over all other aggregation strategies.
[…]
Commonsense FedAvg 74.89
LLaMA-2 7B Mean_ABM 75.15 +0.25
GeoMed_ABM 76.55 +1.66
Arithmetic FedAvg 24.87
LLaMA-2 7B Mean_ABM 25.12 +0.25
GeoMed_ABM 26.09 +1.22
GLUE FedAvg 93.57
RoBERTa Mean_ABM 94.10 +0.53
GeoMed_ABM 94.91 +1.34
@@
To solve this optimization efficiently, we employ Weiszfeld’s algorithm, an iterative method known to converge under mild conditions. Details of the algorithm are provided in Table 1, Figure 5, and Appendix D.4. By avoiding direct averaging and incorporating semantically meaningful low-rank intervention updates through robust aggregation, ABM enables each client to benefit from the knowledge of others without sacrificing local personalization. This approach enhances stability and generalization across non-i.i.d. and task-heterogeneous FL environments. Geometric median has complexity of O(T·d), with T being the number of iterations and d the number of parameters.
3.4 Local Model Update with ABM Aggregation using Test-Time Computing
In FedReFT, local interventions capture client-specific semantics, while ABM aggregation conveys shared global knowledge. However, naively averaging representation-level updates can risk disrupting semantic consistency across clients, potentially leading to misaligned feature spaces. Moreover, uniform aggregation overlooks client heterogeneity, limiting adaptability in diverse edge environments.
These challenges highlight the need for a mechanism that can dynamically balance personalization and global alignment. Without such adaptation, clients may either overfit to their own data or lose semantic fidelity when forced into uniform global updates. FedReFT introduces an adaptive mixing strategy inspired by Test-Time Computing (TTC) wang2021tent, which learns client-specific coefficients to combine local and ABM aggregated intervention parameters. Unlike conventional fine-tuning that updates model weights, TTC performs lightweight optimization during inference, using a small, controlled compute budget to iteratively refine predictions and learn adaptive mixing parameters. Here, unconstrained logits $\beta\in\mathbb{R}$ define mixing weights $\alpha=\sigma(\beta)\in(0,1)$ via the sigmoid $\sigma(\cdot)$. For any intervention tensor $\mathit{X}\in{\mathit{W},\mathit{R},\mathit{B}}$, the mixed parameter is:
$\mathit{X}_{\text{mixed}}(\alpha)\;=\;\alpha\,\mathit{X}^{\text{local}}+(1-\alpha)\,\mathit{X}^{\text{ABM}}.$ (8)
TTC runs for $S$ steps over $B$ calibration batches of clients’ test data. At each step, it injects $\mathit{X}_{\text{mixed}}(\alpha)$ into the model interventions, computes a forward pass to evaluate the test loss $\mathcal{L}_{\text{test}}$, backpropagates gradients $\nabla_{\beta}\mathcal{L}$ while keeping $\mathit{X}^{\text{local}}$ and $\mathit{X}^{\text{ABM}}$ frozen, and then updates $\beta\leftarrow\beta-\eta\nabla_{\beta}\mathcal{L}$. The gradient $\nabla_{\beta}\mathcal{L}=\nabla_{\alpha}\mathcal{L}\,\sigma(\beta)\big(1-\sigma(\beta)\big)$ directly indicates whether increasing $\alpha$ reduces the test loss: negative gradients push $\beta$ higher (increasing $\alpha$), while positive gradients push it lower. Over iterations, $\beta$ converges to:
$\beta^{*}\;=\;\arg\min_{\beta}\;\mathcal{L}_{\text{test}}\!\big(\mathit{X}_{\text{mixed}}(\sigma(\beta))\big)\;+$ (9)
$\lambda_{1}\mathcal{H}(\alpha)\;+\;\lambda_{2}\mathcal{C}(\alpha)\;+\;\lambda_{3}\mathcal{D}(\alpha),$
where $\mathcal{H}$ (entropy) discourages collapse, $\mathcal{C}$ (consistency) promotes stable mixing across keys, and $\mathcal{D}$ (diversity) avoids extreme $\alpha$ saturation.
Final Mixed Interventions. The optimal $\alpha^{*}=\sigma(\beta^{*})$ is used to mix the local and ABM intervention parameters:
$\mathit{X}_{k}^{\text{new}}$ $=\alpha^{*}\mathit{X}_{k}^{\text{local}}+(1-\alpha^{*})\mathit{X}_{k}^{\text{ABM}},$ (10)
$\mathit{X}$ $\in\{\mathit{W},\mathit{R},\mathit{B}\}.$
Before the next local training with $\mathit{R}_{k}^{\text{new}}$, we apply an orthogonal transformation to $\mathit{R}_{k}^{\text{new}}$ to preserve the original property of $R$. We continue the discussion in Appendix D.3 and Table LABEL:tbl:baseline_first_commmonsense_ablation, LABEL:tbl:glue_task_ablation, 14, and LABEL:tbl:math_setup_3_baseline_ablation show that the performance of TTC-based adaptive mixing is over the balanced mixing.
Computational Overhead. The TTC process costs $S\times B$ forward/backward passes per client. With $N$ clients, $K$ intervention parameters, batch size $B$, $L$ layers, and hidden width $d_{h}$, an approximate FLOPs is $\mathcal{O}\!\left(S\cdot M\cdot N\cdot K\cdot B\cdot L\cdot d_{h}^{2}\right)$. Amortized over $T$ communication rounds and $|\mathcal{D}_{\text{test}}|$ inference batches, the effective per-inference overhead scales as $\frac{S\cdot B}{T\cdot|\mathcal{D}_{\text{test}}|}$. Here, small S=50S{=}50 and B=8B{=}8 keep the overhead modest while still improving robustness under shift.
@@ 3.4 Local Model Update with ABM Aggregation using Test-Time Computing
Table 2: Performance of LLaMA-3.2 3B across five commonsense reasoning tasks with Mixed-Task (MT) setup, where clients train on heterogeneous task mixtures to promote generalizable representations. *Performance results of all baseline methods and the experimental setup are taken from singhal2025fed. Trainable Parameter (TP) Efficiency Rank 8 quantifies the relative parameter efficiency of FedReFT(R8), indicating how many times fewer trainable parameters it requires compared to baseline methods.
[…]
Table 3: Performance of Federated fine-tuning of Llama-3.2 3B across eight commonsense reasoning datasets in a highly data-heterogeneous setting, which is denoted as Distinct Task (DT). *Performance results of all baseline methods and the experimental setup are taken from singhal2025fed. Trainable Parameter (TP) Efficiency Rank quantifies the relative parameter efficiency of FedReFT(R8). Best results are in bold.
[…]
FedIT* 32 48.63 60.89 78.22 69.92 73.18 67.78 81.21 67.04 66.91 70.80
FFA-LoRA* 32 24.31 60.73 76.91 65.37 68.61 61.89 79.41 62.92 63.12 67.17
FedEx-LoRA* 32 243.15 62.55 79.36 71.41 71.78 72.45 82.69 67.80 70.25 73.13
FLoRA* 32 243.15 62.55 79.36 71.41 71.78 72.45 82.69 67.80 70.25 73.13
Fed-SB* 200 7.85 63.28 80.34 73.56 82.07 76.01 84.01 69.02 72.46 75.21
FedReFT(ours) 8 2.75 63.84 78.01 77.53 89.74 72.30 84.57 69.31 77.00 76.41
16 5.50 64.67 81.88 78.06 89.88 77.82 86.11 69.37 76.20 78.00
Table 4: Performance comparison across arithmetic reasoning tasks with the Distinct Task (DT) and Mixed Task (MT) setup using different model sizes. We report results under adaptive mixing with Test-Time Computing, which dynamically balances local and global interventions at inference. Avg represents the average of the clients’ accuracy.
@@
[…]
LLaMA 7B 26.12 26.71 49.94 34.26 20.86 15.60 28.22 22.23
LLaMA-2 7B 30.96 33.31 59.51 41.93 22.05 23.50 32.74 26.09
LLaMA-3 8B 35.36 49.41 75.25 53.34 33.45 51.28 73.51 52.75
@@ 4 Experimental Validation
To evaluate FedReFT, we conduct extensive experiments on three different NLP benchmarks covering over 12 datasets. Our objective is to present a comprehensive assessment of how this approach performs in various NLP tasks. We experiment with both masked and autoregressive language models, including RoBERTa-large, TinyLLaMA-1B, LLaMA 7B, LLaMA-2 7B and 13B, LLaMA-3.2B and LLaMA-3 8B, across multiple settings and scales. Our comparisons include state-of-the-art baselines, such as LoRAhu2021lora, FedIT zhang2024towards, FFA-LoRAsun2024improving, FedDPA-LoRAlong2024dual, FedSA-LoRAguo2024selective, Fed-SB singhal2025fed and FLoRAwang2024flora focusing on both parameter efficiency and performance trade-offs. We align the experimental setup configurations with the baseline papers to ensure fair comparisons. To optimize memory usage, we load all base language models with torch.bfloat16 precision. The results are averaged over two runs to report the mean performance.
Hyperparameter Configuration. In the experiments, we determine the number of interventions to learn and the specific layers and input positions where they are applied. Interventions are inserted at a fixed number of layers ($L$) and at the prefix ($p$) and suffix ($s$) positions of the input prompt. We narrow the hyperparameter search space for FL in Appendix B by adopting the configuration used in the centralized ReFTWu2024 paper.
Task Distribution Rationale. We design two experimental setups for commonsense and arithmetic reasoning tasks to study how global representations converge under diverse task distributions. In the Mixed-Task (MT) setup, each client trains on a subset of a combined reasoning dataset but is evaluated on a single task, encouraging generalized, transferable representations through ABM aggregation. This reflects collaborative learning across varied yet related tasks. In the Distinct Task (DT) setup, each client trains on a unique reasoning task, enabling personalized fine-tuning while still leveraging global updates. Despite higher task heterogeneity, this setup maintains stable performance as model capacity increases. Both setups show that FedReFT supports effective generalization in MT and robustness in DT.
@@ 4.1 Commonsense Reasoning
We evaluate global representation generation on eight commonsense reasoning tasks using the Commonsense170K dataset inspired by singhal2025fed; Wu2024. We use the same hyperparameter of  singhal2025fed and tune the intervention parameter in the Appendix B.1. This helps us tune important hyperparameters efficiently and also test their robustness across multiple commonsense reasoning tasks.
Datasets. For the first setup, as a Mixed-Task (MT) design, we split the commonsense reasoning tasks dataset Commonsense170K hu2023llm among clients and use them for fine-tuning. Each client evaluates one of the commonsense reasoning tasks. BoolQ, PIQA, SIQA, HellaSwag, and WinoGrande. For the second setup as Distinct Task (DT) design, each client fine-tunes on only one of these five commonsense reasoning tasks and is evaluated using the same task. We adopt the prompt template from Hu et al. hu2023llm. The Distinct Task (DT) experiment is described in the Appendix 16, as no baseline was found in this setting.
Table 5: Performance comparison across GLUE Tasks on RoBERTa model for $C=3$, FedReFT use rank 1. *Performance results of all baseline methods are taken from guo2024selective and use LoRA rank 8. FedReFT achieves higher accuracy over all the baselines in FL settings. (TP) Efficiency quantifies the relative trainable parameter efficiency of FedReFT, indicating how many times fewer parameters it requires compared to baselines.
@@
[…]
@@ 4.2 Arithmetic Reasoning
Datasets. In the first setting following Mixed-Task (MT), we split the arithmetic reasoning dataset, MATH10K hu2023llm, which includes four tasks with chain-of-thought solutions generated by a language model. Each client reports performance using test set one of three tasks: AQuA, SVAMP, and MAWPS. In the second setting following Distinct-Task (DT), each client is assigned one arithmetic reasoning task for both fine-tuning and evaluation. In the third setting, following guo2024selective; kuang2024federatedscope, we split the dataset GSM8K into 3 clients under an IID distribution.
Results. In Table LABEL:tbl:math_setup_3_baseline, FedReFT achieves the highest accuracy among all methods while requiring substantially fewer trainable parameters, underscoring its efficiency and superior performance. In Table 4, the DT setup represents task heterogeneity, enabling clients to learn personalized, task-specific representations while benefiting from global aggregation. Consequently, the DT setup achieves higher performance. In contrast, the Mixed-Task (MT) setup trains clients on heterogeneous task mixtures to promote globally generalizable representations, but this blending can reduce performance on specific evaluation tasks due to representation misalignment and conflicting objectives.
[…]
@@ 4.3 Natural Language Understanding
We evaluate the effectiveness of FedReFT in learning generalizable representations for Natural Language Understanding (NLU) using the GLUE benchmark wang2018glue. The objective is to fine-tune NLU to learn global representations that capture task-level semantics. By aligning intermediate representations for downstream classification performance. This setup allows us to test whether lightweight intervention tuning can align representations across clients within a single NLU task.
[…]
@@ 4.4 Ablation Studies
We conduct ablation studies in Appendix D to further evaluate the effectiveness of FedReFT, focusing on the impact of geometric-median-based All-But-Me aggregation, intervention parameter sharing strategies, and local model update approaches using balanced and TTC-based adaptive mixing
@@ 5 Conclusion
In this work, we bridge the gap between Representation Fine-Tuning (ReFT) and Federated Learning by introducing FedReFT, a unified framework that enables personalized and parameter-efficient federated representation learning. FedReFT employs the All-But-Me aggregation strategy to mitigate semantic misalignment caused by naive averaging, enabling clients to adapt using a robust average of others’ interventions. Additionally, an adaptive mixing mechanism inspired by Test-Time Computing dynamically balances local and global representations, enhancing robustness under heterogeneous conditions. Extensive experiments demonstrate that FedReFT consistently improves convergence, generalization, and parameter efficiency across diverse FL settings.
@@
Data and Model Usage
We use publicly available models including LLaMA-1.1B, LLaMA-2 (7B, 13B), LLaMA-3 8B, LLaMA-3.2 3B and RoBERTa-large. LLaMA-2 and LLaMA-3 models are licensed under Meta’s community license permitting commercial use. RoBERTa-large is under the MIT License, and TinyLLaMA use Apache 2.0, while the original LLaMA-1 7B is for non-commercial research only. We will release code and configurations under an open-source license with usage documentation to support reproducibility and responsible use.
We employ publicly available datasets across commonsense and arithmetic reasoning tasks, each released under open-source licenses. For commonsense reasoning, BoolQ is under CC BY-SA 3.0, PIQA under Apache 2.0, SIQA and WinoGrande under CC BY 4.0, HellaSwag under MIT, ARC under CC BY-SA 4.0, and OBQA under CC BY 4.0. For arithmetic reasoning, AddSub, AQuA, MAWPS, and MultiArith are under Apache 2.0, GSM8K and SVAMP under MIT, and SingleEq under CC BY 4.0. For natural language understanding, GLUE consists of multiple datasets, each with its own license, allowing for research use and redistribution.
Environmental Impact
[…]
Societal Impacts
Our method FedReFT adapts ReFT for Federated Learning, enabling efficient model personalization with minimal computational overhead. This promotes broader accessibility of large language models on edge devices, including in low-resource or privacy-sensitive environments. While improving inclusivity and deployment scalability, care must be taken to mitigate potential misuse or bias propagation across decentralized systems.
Bias and Fairness
Our approach FedReFT considers the potential for bias introduced by non-IID client data in Federated Learning. While we do not explicitly optimize for fairness, we acknowledge that imbalanced participation or data diversity may lead to uneven model performance. Future work should explore fairness-aware objectives to mitigate such disparities across clients and demographic groups.
Responsible Deployment
To support responsible use, we include clear documentation outlining the intended use cases of our framework and advise against applying it in safety-critical settings without thorough validation. We encourage users to follow ethical standards, such as the ACL Code of Ethics, when deploying our method. Our released code comes with usage instructions to promote safe adoption and reduce the risk of misuse. This work is licensed under CC BY 4.0, allowing reuse and adaptation, even commercially, with proper attribution.
AI Assistants in Research Writing
We used AI assistants to support writing and code refinement during the preparation of this paper. All AI-generated content was reviewed and verified by the authors.